OpenStudy (anonymous):

Is it wrong that the solution of the worked example in session 69?

OpenStudy (anonymous):

can't you post the example ?

OpenStudy (chalisc):

The answer is incorrect. Upon closer inspection I believe that the solver didn't get the integration of cos^3 (2x) correct. Answer should be: \[\dfrac{1}{192}\left( \sin (6x) - 3 \sin (4x) - 3 \sin (2x) + 12x\right) + C\] Note: Answers will vary depending on what trig identities were utilized.